The length of a rectangle is 2m more than the width. The area of the rectangle is 48m squared. What is the length and width?
1
Expert's answer
2020-04-15T17:32:14-0400

Let the width be equal to x, the length be equal to x+2. We know that area of rectangle is a product of length and width. So, we have S(area) =width *length. Now substitute the values, 48=x*(x+2) and solve the equation.

x2+2x-48=0

x2+2x+1-49=0

x+1)2-72=0

(x+1-7)(x+1+7)=0

(x-6)(x+8)=0

x-6=0 or x+8=0

X=6     x=-8, but a negative value is not possible.

Thus, the width is 6m and the length is 8m.